OpenStudy (anonymous):

use Laplace transforms to solve the system: y(dot) +6y=x(dot), 3x-x(dot)=2y(dot); x(0)=2, y(0)=3

OpenStudy (anonymous):

What do dot represent?

OpenStudy (anonymous):

the derivatives over time so x(dot)=dx/dt and y(dot)=dy/dt

OpenStudy (anonymous):

y'+6y=x' 3x-x'=2y' sY(s)-3 + 6 Y(s)= s X(s)-2 3X(s)-(S X(s)-2)= 2 sY(s)- 3 x(0)=2 y(0)=3

OpenStudy (anonymous):

Well, we got two equation , two unknown

OpenStudy (anonymous):

Have to solve for X(s) and Y(s)

OpenStudy (anonymous):

i cant seem to get an answer keep on getting stuck.. Could you help me out?

OpenStudy (anonymous):

good you are still here

OpenStudy (anonymous):

Check my algebra Y (s) (s + 6) - s X (s) = 1 Y (s) (-2 s) + (3 - s) X (s) = -5

OpenStudy (anonymous):

\[\left( \begin{array}{cc} s+6 & -s \\ -2 s & 3-s \end{array} \right)\left( \begin{array}{c} Y(s) \\ X(s) \end{array} \right)=\left( \begin{array}{c} 1 \\ -5 \end{array} \right)\]

OpenStudy (anonymous):

A x =B x= A^-1 B

OpenStudy (anonymous):

made a stupid error there.. wrote -1 instead of -5

OpenStudy (anonymous):

When solved Y(s)=\[\frac{3-s}{18-3 s-3 s^2}-\frac{5 s}{18-3 s-3 s^2}\]

OpenStudy (anonymous):

X(s)=\[\frac{2 s}{18-3 s-3 s^2}-\frac{5 (6+s)}{18-3 s-3 s^2}\]

OpenStudy (anonymous):

y(t)=\[\frac{1}{5} e^{-3 t} \left(7+3 e^{5 t}\right)\] x(t)=\[\frac{1}{5} e^{-3 t} \left(-7+12 e^{5 t}\right)\]

OpenStudy (anonymous):

Let's check \[y'[t]+6y[t]=\frac{3}{5} e^{-3 t} \left(7+8 e^{5 t}\right)\] \[x'[t]=\frac{3}{5} e^{-3 t} \left(7+8 e^{5 t}\right)\] They Match!!!!, We are good

OpenStudy (anonymous):

wow you are simply amazing :)